OpenStudy (anonymous):

Solve for x: 3(x + 1) = −2(x − 1) − 4 A)1 B)–1 C)–5 D)–25

OpenStudy (anonymous):

3(x+1)=-2(x-1)-4 3x+3=-2x+2-4 3x+3=-2x-2 3x+2x=-3-2 5x=-5 x=-5/5=-1

OpenStudy (anonymous):

so your answer is B

OpenStudy (anonymous):

can you message me, I need help with a couple of more. like 3. But I I don't want to post. /:

OpenStudy (anonymous):

okay, well do you get the one i just worked?

OpenStudy (anonymous):

Yes. I do.